1) Find the tangent line to the graph given by x^2(x^2+y^2)=y^2 at the points(√2/2, √2/2)


1
Expert's answer
2022-02-28T11:39:45-0500

Differentiate both sides with respect to "x"


"(x^2(x^2+y^2))'=(y^2)'"

Use the Chain Rule


"4x^3+2xy^2+2x^2yy'=2yy'"

Solve for "y'"


"y'=\dfrac{4x^3+2xy^2}{2y(1-x^2)}"

Point "(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})"


"y'(\dfrac{\sqrt{2}}{2})=\dfrac{4(\dfrac{\sqrt{2}}{2})^3+2(\dfrac{\sqrt{2}}{2})(\dfrac{\sqrt{2}}{2})^2}{2(\dfrac{\sqrt{2}}{2})(1-(\dfrac{\sqrt{2}}{2})^2)}=2"

The tangent line to the graph 


"y-\dfrac{\sqrt{2}}{2}=2(x-\dfrac{\sqrt{2}}{2})"

"y=2x-\dfrac{\sqrt{2}}{2}"